# Soft Theorems in Chern-Simons Matter Theories

### v3

@@ 3 Chern-Simons Vertices in D=4+1
Now that we have the subleading soft theorems we can move to the Chern-Simons correction. For this we need to derive the tree level Chern-Simons vertices. Let us look at the case of abelian gauge field first and then the non-abelian.
3.1 Abelian Case - Chern-Simons QED
@@
$\mathcal{S}_{CS}$ $=\kappa\int\mathbb{A}\left(d\mathbb{A}\right)^{2}=\kappa\int d^{4+1}x\,\epsilon^{\mu\nu\rho\sigma\delta}A_{\mu}\partial_{\nu}A_{\rho}\partial_{\sigma}A_{\delta}.$ (30)
@@
$i\mathcal{M}=$ $i\kappa\epsilon^{\mu\nu\rho\sigma\delta}\left[\varepsilon_{\mu}\left(k_{1}\right)ik_{2\nu}\varepsilon_{\rho}\left(k_{2}\right)ik_{3\sigma}\varepsilon_{\delta}\left(k_{3}\right)+\varepsilon_{\mu}\left(k_{1}\right)ik_{3\nu}\varepsilon_{\rho}\left(k_{3}\right)ik_{2\sigma}\varepsilon_{\delta}\left(k_{2}\right)\right$.
$+\varepsilon_{\mu}\left(k_{2}\right)ik_{3\nu}\varepsilon_{\rho}\left(k_{3}\right)ik_{1\sigma}\varepsilon_{\delta}\left(k_{1}\right)+\varepsilon_{\mu}\left(k_{2}\right)ik_{1\nu}\varepsilon_{\rho}\left(k_{1}\right)ik_{3\sigma}\varepsilon_{\delta}\left(k_{3}\right)$
$\left.+\varepsilon_{\mu}\left(k_{3}\right)ik_{1\nu}\varepsilon_{\rho}\left(k_{1}\right)ik_{2\sigma}\varepsilon_{\delta}\left(k_{2}\right)+\varepsilon_{\mu}\left(k_{3}\right)ik_{2\nu}\varepsilon_{\rho}\left(k_{2}\right)ik_{1\sigma}\varepsilon_{\delta}\left(k_{1}\right)\right]$
$=$ $-2i\kappa\epsilon^{\mu\nu\rho\sigma\delta}\left[k_{1\nu}k_{2\sigma}+k_{2\nu}k_{3\sigma}+k_{3\nu}k_{1\sigma}\right]\varepsilon_{\mu}\left(k_{1}\right)\varepsilon_{\rho}\left(k_{2}\right)\varepsilon_{\delta}\left(k_{3}\right).$ (31)
@@
$\Gamma_{CS}^{\mu\rho\delta}\left(k_{1},k_{2},k_{3}\right)=-2i\kappa\epsilon^{\mu\nu\rho\sigma\delta}\left[k_{1\nu}k_{2\sigma}+k_{2\nu}k_{3\sigma}+k_{3\nu}k_{1\sigma}\right]=-6i\kappa\epsilon^{\mu\nu\rho\sigma\delta}k_{1\nu}k_{2\sigma}.$ (32)
@@
3.2 Non-Abelian Case - Chern-Simons QCD
@@
$S_{CS}=\kappa\int Tr\left[\mathbb{A}\left(d\mathbb{A}\right)^{2}+c_{4}\mathbb{A}^{3}d\mathbb{A}+c_{5}\mathbb{A}^{5}\right].$ (33)
@@
Using $\mathcal{S}_{CS}$, the 3-point vertex is given by,
$\Gamma_{CS;abc}^{\left(3\right)\mu\rho\delta}\left(k_{1},k_{2},k_{3}\right)=-\frac{3i}{2}\kappa\epsilon^{\mu\nu\rho\sigma\delta}d_{abc}k_{1\nu}k_{2\sigma},$ (34)
where $\frac{i}{2}d^{abc}=Tr\left[T^{a}\left\{T^{b},T^{c}\right\}\right]$. Similarly we can obtain the 4-point vertex to be,
$\Gamma_{CS;abcd}^{\left(4\right)\mu\nu\rho\delta}\left(k_{1},k_{2},k_{3},k_{4}\right)$
$\,\,=-c_{4}\kappa\epsilon^{\mu\nu\rho\sigma\delta}\left[k_{1\sigma}\left\{f_{dce}Tr\left(T_{b}T_{e}T_{a}\right)+f_{bde}Tr\left(T_{c}T_{e}T_{a}\right)+f_{cbe}Tr\left(T_{d}T_{e}T_{a}\right)\right\}\right$.
$\,\,\,\,\,\,\,+k_{2\sigma}\left\{f_{cde}Tr\left(T_{a}T_{e}T_{b}\right)+f_{dae}Tr\left(T_{c}T_{e}T_{b}\right)+f_{ace}Tr\left(T_{d}T_{e}T_{b}\right)\right\}$
$\,\,\,\,\,\,\,+k_{3\sigma}\left\{f_{dbe}Tr\left(T_{a}T_{e}T_{c}\right)+f_{ade}Tr\left(T_{b}T_{e}T_{c}\right)+f_{bae}Tr\left(T_{d}T_{e}T_{c}\right)\right\}$
$\,\,\,\,\,\,\,\left.+k_{4\sigma}\left\{f_{bce}Tr\left(T_{a}T_{e}T_{d}\right)+f_{cae}Tr\left(T_{b}T_{e}T_{d}\right)+f_{abe}Tr\left(T_{c}T_{e}T_{d}\right)\right\}\right].$ (35)
@@
$\Gamma_{CS;abcd}^{\left(4\right)\mu\nu\rho\delta}\left(k_{1},k_{2},k_{3},k_{4}\right)$
$\,\,=-\frac{c_{4}}{4}i\kappa\epsilon^{\mu\nu\rho\sigma\delta}\left[\left(k_{1\sigma}-k_{2\sigma}\right)f_{dce}d_{bea}+\left(k_{1\sigma}-k_{3\sigma}\right)f_{bde}d_{cea}+\left(k_{1\sigma}-k_{4\sigma}\right)f_{cbe}d_{dea}\right$.
@@
We can also calculate the 5-point vertex $\Gamma_{CS}^{\left(5\right)}$ but for our purpose it will suffice to just note that it will have no momentum dependence at tree level.
@@ 4 Chern-Simons Correction to Single Soft Photon Theorem in QED
Now we add the Chern-Simons term to the action and see how it modifies the soft photon theorem. We start with considering a single soft photon in the external state. Then the relevant diagrams are the ones of type $A1$ and $A2$. From (32), we see that the Chern-Simons vertex modifies $\Gamma^{\left(3\right)}$ only when the external hard particle is the massless $U\left(1\right)$ gauge boson. In that case we have,
$\delta\Gamma^{\left(3\right)\alpha\beta}\left(e,k;p,-p-k\right)=$ $e_{\mu}\Gamma_{CS}^{\mu\alpha\beta}\left(k,-p-k,p\right)$
$=$ $6i\kappa\epsilon^{\mu\nu\alpha\sigma\beta}e_{\mu}k_{\nu}p_{\sigma},$ (37)
@@
$K_{photon}^{\mu\nu}\left(p\right)=-p^{2}\eta^{\mu\nu},$
$\Xi_{\mu\nu}^{photon}\left(p\right)=-i\eta_{\mu\nu}.$
@@
$\delta A1=$ $\varepsilon^{T}\left(p_{i}\right)\delta\Gamma^{\left(3\right)}\left(e,k;p_{i},-p_{i}-k\right)\frac{\Xi_{i}^{photon}\left(-p_{i}-k\right)}{2p_{i}.k}G_{n}^{\left(i\right)}\left(p_{i}+k\right)$
$=$ $3i\kappa\frac{1}{p_{i}.k}\varepsilon_{i;\alpha}\epsilon^{\mu\nu\alpha\sigma\beta}e_{\mu}k_{\nu}p_{i\sigma}\Xi_{\beta\gamma}^{photon}\left(-p_{i}\right)G_{n}^{\left(i\right);\gamma}\left(p_{i}\right)$
@@
$+\sum_{i}\frac{1}{p_{i}.k}\frac{i}{4}\left(e_{\mu}k_{\nu}-e_{\nu}k_{\mu}\right)\varepsilon_{i}^{T}\frac{\partial K_{i}\left(-p_{i}\right)}{\partial p_{i\nu}}\frac{\partial\Xi_{i}\left(-p_{i}\right)}{\partial p_{i\mu}}Q_{i}^{T}G_{n}^{\left(i\right)}\left(p_{i}\right)$
@@ 5 Chern-Simons Correction to Single Soft Gluon Theorem in QCD
$\delta\Gamma_{ab}^{\left(3\right)\alpha\beta}\left(e,k;p,-p-k\right)=$ $e_{c\mu}\Gamma_{CS;cab}^{\mu\alpha\beta}\left(k,-p-k,p\right)$
$=$ $\frac{3i\kappa}{2}\epsilon^{\mu\nu\alpha\sigma\beta}d_{cab}e_{c\mu}k_{\nu}p_{\sigma},$ (39)
@@
$K_{ab}^{gluon;\mu\nu}\left(p\right)=-p^{2}\eta^{\mu\nu}\delta_{ab},$ (40)
$\Xi_{\mu\nu ab}^{gluon}\left(p\right)=-i\eta_{\mu\nu}\delta_{ab}.$ (41)
@@
$=$ $\frac{3i\kappa}{4}\frac{1}{p_{i}.k}\varepsilon_{i;a\alpha}\epsilon^{\mu\nu\alpha\sigma\beta}d_{cab}e_{c\mu}k_{\nu}p_{i\sigma}\Xi_{\beta\gamma bd}^{gluon}\left(-p_{i}\right)G_{n;d}^{\left(i\right);\gamma}\left(p_{i}\right)$
@@
$-i\sum_{i}\varepsilon_{i}^{T}g_{i}\frac{e_{a\mu}k_{\nu}}{p_{i}.k}T_{i;a}\left(\hat{J}_{\left(i\right)}^{\mu\nu}G_{n}^{\left(i\right)}\right)\left(p_{i}\right)$
@@ 6 Chern-Simons Correction to Soft Theorems with Multiple Simultaneous Soft Emissions
$\Gamma_{CS}^{\left(3\right)}\sim k\Gamma_{gauge}^{\left(3\right)},$
$\Gamma_{CS}^{\left(4\right)}\sim k\Gamma_{gauge}^{\left(4\right)}.$
This implies that there are no diagrams involving these vertices that contribute to the leading soft theorem. This can be seen by realizing that if there were any such diagrams we could’ve replaced the Chern-Simons vertices with their pure gauge counterparts and would have gotten an even more dominant contribution. This also implies that the Chern-Simons corrections to subleading soft theorems come largely from adding the Chern-Simons correction to the gauge vertices in the diagrams contributing to the leading soft theorem, as was seen in the case of single soft emission. The only exception are diagrams that have the 5-point Chern-Simons vertex as they do not have a corresponding more dominant gauge vertex. But we can argue that these also lead to only subleading contributions. For instance, consider figure 3 which contributes to the Chern-Simons correction for the subleading soft theorem with four soft emissions. This is the only diagram containing the 5-point vertex which contributes at the subleading order for this process, i.e. at $\mathcal{O}\left(\nicefrac{{1}}{{k^{3}}}\right)$. It is easy to see that there is no diagram containing the 5-point vertex which is more dominant than this since it maximizes the negative $k$ powers in the propagator while having the 5-point Chern-Simons vertex.
$\Gamma^{(3)}$$G_{n}^{(i)}$$\varepsilon_{i}$$p_{i}$$\varepsilon_{i}$$p_{i}+k_{1}+...k_{4}$$k_{1}+...k_{4}$$\varepsilon_{1}$$p_{1}$$...\varepsilon_{i-1}$$p_{i-1}$$\varepsilon_{i+1}$$p_{i+1}$$\varepsilon_{n}$$p_{n}$$e_{4}$$k_{4}$$e_{1}$$k_{1}$$\Gamma_{CS}^{(5)}$......
@@
Similar argument holds for soft theorems with even more soft emissions. Hence, we can safely conclude that even for multiple soft emissions, the Chern-Simons corrections are subleading to the gauge theory.

### v4

@@ 3 Chern-Simons Vertices in D=4+1
Now that we have the subleading soft theorems we can move to the Chern-Simons correction. For this we need to derive the tree-level Chern-Simons vertices. Let us look at the case of Abelian gauge field first and then the non-Abelian.
3.1 Abelian Case: Chern-Simons QED
@@
$S_{\text{CS}}$ $=\kappa\int\mathbb{A}\left(d\mathbb{A}\right)^{2}=\kappa\int d^{4+1}x\,\epsilon^{\mu\nu\rho\sigma\delta}A_{\mu}\partial_{\nu}A_{\rho}\partial_{\sigma}A_{\delta}.$ (28)
@@
$\mathrm{i}\mathcal{M}=$ $\mathrm{i}\kappa\epsilon^{\mu\nu\rho\sigma\delta}\left[\varepsilon_{\mu}\left(k_{1}\right)\mathrm{i}k_{2\nu}\varepsilon_{\rho}\left(k_{2}\right)\mathrm{i}k_{3\sigma}\varepsilon_{\delta}\left(k_{3}\right)+\varepsilon_{\mu}\left(k_{1}\right)\mathrm{i}k_{3\nu}\varepsilon_{\rho}\left(k_{3}\right)\mathrm{i}k_{2\sigma}\varepsilon_{\delta}\left(k_{2}\right)\right$.
$+\varepsilon_{\mu}\left(k_{2}\right)\mathrm{i}k_{3\nu}\varepsilon_{\rho}\left(k_{3}\right)\mathrm{i}k_{1\sigma}\varepsilon_{\delta}\left(k_{1}\right)+\varepsilon_{\mu}\left(k_{2}\right)\mathrm{i}k_{1\nu}\varepsilon_{\rho}\left(k_{1}\right)\mathrm{i}k_{3\sigma}\varepsilon_{\delta}\left(k_{3}\right)$
$\left.+\varepsilon_{\mu}\left(k_{3}\right)\mathrm{i}k_{1\nu}\varepsilon_{\rho}\left(k_{1}\right)\mathrm{i}k_{2\sigma}\varepsilon_{\delta}\left(k_{2}\right)+\varepsilon_{\mu}\left(k_{3}\right)\mathrm{i}k_{2\nu}\varepsilon_{\rho}\left(k_{2}\right)\mathrm{i}k_{1\sigma}\varepsilon_{\delta}\left(k_{1}\right)\right]$
$=$ $-2\mathrm{i}\,\kappa\epsilon^{\mu\nu\rho\sigma\delta}\left[k_{1\nu}k_{2\sigma}+k_{2\nu}k_{3\sigma}+k_{3\nu}k_{1\sigma}\right]\varepsilon_{\mu}\left(k_{1}\right)\varepsilon_{\rho}\left(k_{2}\right)\varepsilon_{\delta}\left(k_{3}\right).$ (29)
@@
$\Gamma_{\text{CS}}^{\mu\rho\delta}\left(k_{1},k_{2},k_{3}\right)=-2\mathrm{i}\,\kappa\epsilon^{\mu\nu\rho\sigma\delta}\left[k_{1\nu}k_{2\sigma}+k_{2\nu}k_{3\sigma}+k_{3\nu}k_{1\sigma}\right]=-6\mathrm{i}\,\kappa\epsilon^{\mu\nu\rho\sigma\delta}k_{1\nu}k_{2\sigma}.$ (30)
@@
3.2 Non-Abelian Case: Chern-Simons QCD
@@
$S_{\text{CS}}=\kappa\int\text{Tr}\left[\mathbb{A}\left(d\mathbb{A}\right)^{2}+c_{4}\mathbb{A}^{3}d\mathbb{A}+c_{5}\mathbb{A}^{5}\right].$ (31)
@@
Using $S_{\text{CS}}$, the 3-point vertex is given by,
$\Gamma_{\text{CS};abc}^{\left(3\right)\mu\rho\delta}\left(k_{1},k_{2},k_{3}\right)=-\frac{3\mathrm{i}}{2}\kappa\epsilon^{\mu\nu\rho\sigma\delta}d_{abc}k_{1\nu}k_{2\sigma},$ (32)
where $\frac{\mathrm{i}}{2}d^{abc}=\text{Tr}\left[T^{a}\left\{T^{b},T^{c}\right\}\right]$. Similarly we can obtain the 4-point vertex to be,
$\Gamma_{\text{CS};abcd}^{\left(4\right)\mu\nu\rho\delta}\left(k_{1},k_{2},k_{3},k_{4}\right)$
$\,\,=-c_{4}\kappa\epsilon^{\mu\nu\rho\sigma\delta}\left[k_{1\sigma}\left\{f_{dce}\text{Tr}\left(T_{b}T_{e}T_{a}\right)+f_{bde}\text{Tr}\left(T_{c}T_{e}T_{a}\right)+f_{cbe}\text{Tr}\left(T_{d}T_{e}T_{a}\right)\right\}\right$.
$\,\,\,\,\,\,\,+k_{2\sigma}\left\{f_{cde}\text{Tr}\left(T_{a}T_{e}T_{b}\right)+f_{dae}\text{Tr}\left(T_{c}T_{e}T_{b}\right)+f_{ace}\text{Tr}\left(T_{d}T_{e}T_{b}\right)\right\}$
$\,\,\,\,\,\,\,+k_{3\sigma}\left\{f_{dbe}\text{Tr}\left(T_{a}T_{e}T_{c}\right)+f_{ade}\text{Tr}\left(T_{b}T_{e}T_{c}\right)+f_{bae}\text{Tr}\left(T_{d}T_{e}T_{c}\right)\right\}$
$\,\,\,\,\,\,\,\left.+k_{4\sigma}\left\{f_{bce}\text{Tr}\left(T_{a}T_{e}T_{d}\right)+f_{cae}\text{Tr}\left(T_{b}T_{e}T_{d}\right)+f_{abe}\text{Tr}\left(T_{c}T_{e}T_{d}\right)\right\}\right].$ (33)
@@
$\Gamma_{\text{CS};abcd}^{\left(4\right)\mu\nu\rho\delta}\left(k_{1},k_{2},k_{3},k_{4}\right)$
$\,\,=-\frac{c_{4}}{4}\mathrm{i}\kappa\epsilon^{\mu\nu\rho\sigma\delta}\left[\left(k_{1\sigma}-k_{2\sigma}\right)f_{dce}d_{bea}+\left(k_{1\sigma}-k_{3\sigma}\right)f_{bde}d_{cea}+\left(k_{1\sigma}-k_{4\sigma}\right)f_{cbe}d_{dea}\right$.
@@
We can also calculate the 5-point vertex $\Gamma_{\text{CS}}^{\left(5\right)}$ but for our purpose it will suffice to just note that it will have no momentum dependence at tree-level.
@@ 4 Chern-Simons Correction to Single Soft Photon Theorem in QED
Now we add the Chern-Simons term to the action and see how it modifies the soft photon theorem. We start with considering a single soft photon in the external state. Then the relevant diagrams are the ones of type $A1$ and $A2$. From (30), we see that the Chern-Simons vertex modifies $\Gamma^{\left(3\right)}$ only when the external hard particle is the massless $\text{U}\left(1\right)$ gauge boson. In that case we have,
$\delta\Gamma^{\left(3\right)\alpha\beta}\left(e,k;p,-p-k\right)=$ $e_{\mu}\Gamma_{\text{CS}}^{\mu\alpha\beta}\left(k,-p-k,p\right)$
$=$ $6\mathrm{i}\kappa\epsilon^{\mu\nu\alpha\sigma\beta}e_{\mu}k_{\nu}p_{\sigma},$ (35)
@@
$K_{\text{photon}}^{\mu\nu}\left(p\right)=-p^{2}\eta^{\mu\nu},$
$\Xi_{\mu\nu}^{\text{photon}}\left(p\right)=-\mathrm{i}\eta_{\mu\nu}.$
@@
$\delta A1=$ $\varepsilon^{T}\left(p_{i}\right)\delta\Gamma^{\left(3\right)}\left(e,k;p_{i},-p_{i}-k\right)\frac{\Xi_{i}^{\text{photon}}\left(-p_{i}-k\right)}{2p_{i}.k}G_{n}^{\left(i\right)}\left(p_{i}+k\right)$
$=$ $3\mathrm{i}\kappa\frac{1}{p_{i}.k}\varepsilon_{i;\alpha}\epsilon^{\mu\nu\alpha\sigma\beta}e_{\mu}k_{\nu}p_{i\sigma}\Xi_{\beta\gamma}^{\text{photon}}\left(-p_{i}\right)G_{n}^{\left(i\right);\gamma}\left(p_{i}\right)$
@@
$+\sum_{i}\frac{1}{p_{i}.k}\frac{\mathrm{i}}{4}\left(e_{\mu}k_{\nu}-e_{\nu}k_{\mu}\right)\varepsilon_{i}^{T}\frac{\partial K_{i}\left(-p_{i}\right)}{\partial p_{i\nu}}\frac{\partial\Xi_{i}\left(-p_{i}\right)}{\partial p_{i\mu}}Q_{i}^{T}G_{n}^{\left(i\right)}\left(p_{i}\right)$
@@ 5 Chern-Simons Correction to Single Soft Gluon Theorem in QCD
$\delta\Gamma_{ab}^{\left(3\right)\alpha\beta}\left(e,k;p,-p-k\right)=$ $e_{c\mu}\Gamma_{\text{CS};cab}^{\mu\alpha\beta}\left(k,-p-k,p\right)$
$=$ $\frac{3\mathrm{i}\kappa}{2}\epsilon^{\mu\nu\alpha\sigma\beta}d_{cab}e_{c\mu}k_{\nu}p_{\sigma},$ (37)
@@
$K_{ab}^{\text{gluon};\mu\nu}\left(p\right)=-p^{2}\eta^{\mu\nu}\delta_{ab},$ (38)
$\Xi_{\mu\nu ab}^{\text{gluon}}\left(p\right)=-\mathrm{i}\eta_{\mu\nu}\delta_{ab}.$ (39)
@@
$=$ $\frac{3\mathrm{i}\kappa}{4}\frac{1}{p_{i}.k}\varepsilon_{i;a\alpha}\epsilon^{\mu\nu\alpha\sigma\beta}d_{cab}e_{c\mu}k_{\nu}p_{i\sigma}\Xi_{\beta\gamma bd}^{\text{gluon}}\left(-p_{i}\right)G_{n;d}^{\left(i\right);\gamma}\left(p_{i}\right)$
@@
$-\mathrm{i}\sum_{i}\varepsilon_{i}^{T}g_{i}\frac{e_{a\mu}k_{\nu}}{p_{i}.k}T_{i;a}\left(\hat{J}_{\left(i\right)}^{\mu\nu}G_{n}^{\left(i\right)}\right)\left(p_{i}\right)$
@@ 6 Chern-Simons Correction to Soft Theorems with Multiple Simultaneous Soft Emissions
$\Gamma_{\text{CS}}^{\left(3\right)}\sim k\,\Gamma_{\text{gauge}}^{\left(3\right)},$
$\Gamma_{\text{CS}}^{\left(4\right)}\sim k\,\Gamma_{\text{gauge}}^{\left(4\right)}.$
This implies that there are no diagrams involving these vertices that contribute to the leading soft theorem. This can be seen by realizing that if there were any such diagrams we could have replaced the Chern-Simons vertices with their pure gauge counterparts and would have gotten an even more dominant contribution. This also implies that the Chern-Simons corrections to subleading soft theorems come largely from adding the Chern-Simons correction to the gauge vertices in the diagrams contributing to the leading soft theorem, as was seen in the case of a single soft emission. The only exception are diagrams that have the 5-point Chern-Simons vertex as they do not have a corresponding more dominant gauge vertex. But we can argue that these also lead to only subleading contributions. For instance, consider figure 3 which contributes to the Chern-Simons correction for the subleading soft theorem with four soft emissions. This is the only diagram containing the 5-point vertex which contributes at the subleading order for this process, i.e. at $\mathcal{O}\left(\nicefrac{{1}}{{k^{3}}}\right)$. It is easy to see that there is no diagram containing the 5-point vertex which is more dominant than this since it maximizes the negative $k$ powers in the propagator while having the 5-point Chern-Simons vertex.
$\Gamma^{(3)}$$G_{n}^{(i)}$$\varepsilon_{i}$$p_{i}$$\varepsilon_{i}$$p_{i}+k_{1}+...k_{4}$$k_{1}+...k_{4}$$\varepsilon_{1}$$p_{1}$$...\varepsilon_{i-1}$$p_{i-1}$$\varepsilon_{i+1}$$p_{i+1}$$\varepsilon_{n}$$p_{n}$$e_{4}$$k_{4}$$e_{1}$$k_{1}$$\Gamma_{\text{CS}}^{(5)}$......
@@
Similar argument holds for soft theorems with even more soft emissions. Hence, we can safely conclude that even for multiple soft emissions, the Chern-Simons corrections to soft theorem for tree-level amplitudes are subleading to the gauge theory.
7 Chern-Simons Terms in Higher Dimensions
Finally let us comment on what happens if you add a perturbative Chern-Simons terms to the gauge theory in space-time dimensions other than five. As we stated earlier, in $D=2+1$ the Chern-Simons term is more dominant than the gauge term in the infrared. Therefore, we cannot treat the Chern-Simons term as a perturbation over the gauge theory in three dimensions. In $D>4+1$, the 3-point Chern-Simons vertex is absent [^4] For instance, in $D=6+1$ the lowest vertex is the 4-point vertex which goes as $\sim\int\mathbb{A}d\mathbb{A}^{3}$. Therefore, it is expected that the Chern-Simons term will not correct the subleading soft theorem at the subleading order in $D>4+1$. Therefore, one concludes that the universality of the subleading soft theorem is still maintained in $D>4+1$.